what is the four consecutive numbers of 14
Answer by mathybrain(2)

To find consecutive numbers of a specific number, it is easier to assign each numbers with a letter.
Consecutive numbers=a, specific number=b
You need to first divide b by a.
b/a=14/4
The answer to 14/4 is 3.5
However, because we need to find 4 consecutive numbers, we start to find whole numbers surrounding the answer.
3.5 rounded up is 4 while rounded down is 3.
one less than 3 is 2 while 1 more than 4 is 5.
2+3+4+5=14
2,3,4 and 5 is the four consecutive numbers of 14